OpenStudy (deadshot):

Simplify quantity 12 x minus 96 over quantity x squared minus 4x minus 32 all over quantity 6 x squared minus 6 x minus 12 over 3 x plus 3

OpenStudy (campbell_st):

you question is \[\frac{\frac{12x - 96}{x^2 - 4x - 32}}{\frac{6x^2 -6x -12}{3x + 3}}\] well flip the bottom faction and multiply \[\frac{12x -96}{x^2 -4x -32} \times \frac{3x + 3}{6x^2 - 6x -12}\] but before multiplying... simplify the fractions. \[\frac{12(x - 8)}{(x + 4)(x -8)} \times \frac{3( x + 1)}{6(x-2)(x+1)}\] cancel the common factors then multiply.

OpenStudy (deadshot):

Okay, Thanks!